OpenStudy (anonymous):

What are the solutions of the system? y = x^2 - 3x + 2 y = 4x - 4 a.(1, 0) and (–6, –28) b.(1, 20) and (6, 0) c. (1, 0) and (6, 20) d. no solution

OpenStudy (campbell_st):

equate the functions \[4x - 4 = x^2 - 3x + 2\] simplifying gives \[x^2 - 7x + 6 = 0\] solve the qadratic for x by factorising then substitute into one of the original equations to get y

OpenStudy (anonymous):

both are equal to y, so why not set them equal to each other :)? you can proceed from there, isolating x!

OpenStudy (anonymous):

actually, i was wrong. Quadratic or graph for sure, didn't notice the x^2 part.